# Quantum annealing applications, challenges and limitations for optimisation problems compared to classical solvers

### v3

@@ 1 Introduction
In gate-based QC, the classical, binary bits are replaced by controllable qubits capable of being in non-binary states. Adiabatic quantum annealers, however, function on a different principle. They prepare an ensemble of qubits in the ground state of some initial Hamiltonian, before allowing the system to evolve abiding by the adiabatic theorem [22] to some final state encoding the desired solution. This principle will be further explained in Section 2.1.
@@
This research aims to address this gap by utilising the largest publicly available adiabatic quantum annealers, last updated in August 2023, to solve various optimisation problems, including integer/binary linear programming (BLP) with linear and quadratic constraints, integer/binary quadratic programming (BQP), and mixed-integer linear programming. To assess performance, we systematically compare D-Wave’s Advantage quantum annealers, as described in [40], with the state-of-the-art classical solvers CPLEX [41], Gurobi [42], and IPOPT [43]. Recognising the limitations of fully QC algorithms, we employ D-Wave’s hybrid solvers [44]. This approach seamlessly integrates the strengths of both classical and quantum computation, as explained in more detail in Section 2.5. This study serves as a valuable resource for researchers across various fields dealing with optimisation problems, enabling them to assess the suitability of current D-Wave quantum annealers for their respective applications.
The subsequent sections of the paper are organised as follows: Section 2 provides a brief introduction to the physical and technical background of quantum annealing, encompassing the adiabatic theorem and D-Wave’s architecture in their treatment of constraints and the hybrid computing approach employed. Section 3 presents various case studies for different classes of optimisation problems. Then, Section 4 showcases the application of D-Wave’s hybrid constrained quadratic model (CQM) solver [45] on an energy system problem formulated as Mixed-Integer Linear Programming (MILP).
@@ 2.1 Adiabatic theorem
To perform calculations using a quantum annealer, the physical collection of qubits, hereafter referred to as the system, is prepared in an easy-to-solve initial Hamiltonian $H_{I}$. After the system has been prepared in $H_{I}$, external magnetic fields are applied to change the initial $H_{I}$ into a complex Hamiltonian $H_{F}$. Suppose this process follows the adiabatic theorem described above. In that case, the system will remain in its ground state for the duration of the anneal, and the final state will be the solution of the complex Hamiltonian. As $[\sigma_{x},\sigma_{z}]\neq 0$, $H_{I}$ and $H_{F}$ do not share a common set of eigenstates. As a result, the system must undergo quantum transitions to evolve from one eigenstate to another. If the two Hamiltonians did commute, the eigenstates would remain unchanged throughout the evolution, preventing any quantum dynamics from occurring. The procedure can slowly anneal the system from the quantum Hamiltonian (Equation 5) to the solution of the classical Ising spin system shown in Equation 2. After the annealing process, the collapsed spin states will present a low-energy solution. The annealing quantum Ising Hamiltonian is as follows:
@@
The annealing process begins at $t=0$ and ends at time $t=t_{f}$, where $A(0)\gg B(0)$ and $A(t_{f})\ll B(t_{f})$. At time $t=t_{f}$, the qubits have been de-phased to a classical system and the Pauli matrix can be replaced by the classical $s_{i}=\pm 1$, resulting in Equation 2.
@@ 2.2 D-Wave architecture
The most powerful QPU available is the Pegasus topology [50]. In the Pegasus topology, each qubit is coupled to 15 other qubits, which is a great improvement from the 6 couplings of the previous generation. It consists of a square lattice of unit cells which themselves contain 24 qubits. The Advantage QPU, used for the computations performed in this work, is a 16x16 lattice of unit cells, totalling $24\times 16\times(16-1)=5760$ qubits. Due to the densely packed spatial arrangement of qubits and the physical manufacturing process, it is expected that some qubits may exhibit suboptimal performance. D-Wave estimates that approximately 5% of the qubits in the Pegasus topology may have imperfections. This dense graph structure of the Pegasus systems has allowed for a significant increase in the number of useful qubits, making the Advantage QPU the most promising system to solve the optimisation problems presented in Sections 3.1, 3.4 and 3.3.
@@ 2.3 Constraints
where $b$ is some constant that has to be met. However, when confronted with a linear inequality constraint of the form $P({\bf{x}})=\left(\sum_{i}^{N}a_{i}x_{i}-b\right)\leq 0$, auxiliary slack variables, ${\bf{y}}=\{y_{j}:\forall j\in J\}$ and their corresponding weights $w_{j}$ are used to transform inequality constraints into equality constraints [56] so that Equation 6 can be written as
@@
with coupling values $Q$ and summing over all constraints $k$. With this approach, constrained optimisation problems can be reformulated into a solvable QUBO. We note that other approaches than Lagrangian relaxation have been considered and are actively researched to account for the constraints [57]. Furthermore, the penalty strength $\lambda$ in Equation 8 has to be chosen with care. Contrary to classical computing, there is a far more limited dynamic range of possible pre-factors and coupling strengths that can be applied. Too large values cause a tremendous amount of distortion of the distribution in the ground state. At the same time, too small values will not ensure the satisfaction of the constraint as breaking it could then still result in optimal, unfeasible solutions. The best penalty strength is thereby the lowest possible value that still satisfies the constraints. There is no consensus on how to consistently guarantee this behaviour, and only some guidelines can be found either empirically or algebraically [58, 33].
@@ 3 Case study
This section presents different case studies and their results. The exploration covers various classes of optimisation problems: Integer/Binary Linear Programming (BLP), while also considering the influence of additional constraints of both linear and quadratic nature in Section 3.1, and Integer/Binary Quadratic Programming (BQP) in Section 3.4. To assess the performance of D-Wave’s Advantage system for industrial use, we benchmark the solution quality and computational time of D-Wave’s LeapHybridCQMSolver [45] against results produced by the widely adopted classical solvers CPLEX [41], Gurobi [42], and IPOPT [43] (version 3.14.12). The classical solvers are executed on a dedicated computer cluster equipped with dual 3.5 GHz Intel Xeon Gold 5115 CPUs, each featuring 10 cores and 96 GB of RAM. Noting here that for the time comparison, we measured the total run time as outputted by the cloud service, which includes the time it takes to return the full sample-set of on average 100 sample solutions. From these solutions, we first filter out the feasible set before selecting the solution with the lowest energy value. The time required for these two steps was not included in the computational time measurement for obtaining sufficient solutions. While these steps typically take less than 1s on average, their impact may only be considerable for very short total run times.
@@ 3.1 Integer/Binary Linear Programming (BLP)
where $\mu_{i}$ are random constants between (0,1) and $x_{i}\in\{0,1\}$ are binary variables. This restricts the sum over all binary variables to equal a fixed constant $C$. We first run the solvers with 10,000 to 100,000 binary variables. A comparison of the solutions and the computational time for this range of binary variables is shown in Figure 1(a). Here, we average over 5 runs, each providing a sampleset of approximately 100 solutions from which the best feasible one is selected. As we experienced little deviation in the results we deemed higher statistics not to be necessary. Additionally, as access to quantum computers is still very limited, large statistics are not realisable. IPOPT is designed to handle large-scale nonlinear problems [43] and is not suited for integer variables. As IPOPT approximates the binary variables with continuous ones with very small bounds around zero and one, the derived objective function value with these numbers can score a lower value than what is actually true. We excluded IPOPT’s objective function values in Figure 1(b), however, we kept this solver throughout this work as it shows important results in Section 3.4.
@@
Looking at Figure 1(a), D-Wave stays relatively constant in computation time of about 6s up to 38,000 variables, from whereon it increases more drastically when the number of variables is increased further. Viewing the objective function value, we see that IPOPT behaves similarly to the other solvers for a low number of variables, but starts diverging from the rest as the number of variables increases. It can also be noted that the computational time of IPOPT increases rapidly compared to the other classical solvers. CPLEX and Gurobi start below 1s and gradually increase with an increasing number of variables. D-Wave finds the same optimal solution as Gurobi and CPLEX, resulting in three overlapping lines. The observed sawtooth pattern of the solutions is due to the randomness of the $\mu$ –values as they are newly drawn for each run with a specific seed. When comparing computational time in Figure 1(b), the hybrid solver requires significantly more time as it approaches its scalability limit. This behaviour could be explained by the need for an extremely high amount of partitioning into subproblems as the number of variables increases [35]. Although the optimal solution has been found, D-Wave’s hybrid solvers show no computational advantage.
@@ 3.2 Increasing the complexity by adding more constraints
Next, we focus on determining whether D-Wave could continue to find the optimal solution at its scalability limit when increasing the complexity of the BLP (Equation 9) by adding additional constraints. As the constraints are added as penalty terms, we want to see if the solution quality can be maintained even though the computational time is significantly higher. We consider some arbitrary, although feasible, constraints listed below.
@@
The objective function value and computational time averaged over 5 runs, together with the minimum and maximum deviation as vertical lines, are shown in Figure 2.
@@ 3.3 Integer/Binary Linear Programming with quadratic constraints
In the previous section, all added constraints were linear. The examination, in this section, involves the consideration of a quadratic constraint while keeping the objective function linear. To this end, we added another factor to the constraint from Equation 10 so that it now reads
@@
where $x_{i,j}\in\{0,1\}$, resulting in $N^{2}$ terms. At first, we viewed the solution quality and computational time for an increasing number of variables, shown in Figure 3(a), followed by a stepwise increase in the complexity of the constraint by increasing the value of $C$ in Figure 3(b).
Figure 3: Mean objective function value (top) and computational time (bottom) comparison between two classical solvers and D-Wave’s hybrid solver for the BLP with a single quadratic constraint. In (a) the complexity scales with the number of variables. In (b), the variables are fixed to 500, whereas the complexity scales with increasing $C$ as stated in Equation 10. All results are averaged over 5 runs, and the vertical lines show the minimum and maximum deviation from the mean value. Gurobi’s solve time is significantly larger than all other solvers and is thereby excluded here for visibility.
@@
We experience similar behaviour when stepwise increasing the value of the constraint. D-Wave is again slightly off the optimal solution, however, it stays relatively constant in computational time and solves the problem almost consistently faster than IPOPT. Although the solution is suboptimal, it can be seen in Figure 3(b) that also for quadratic constraints with a linear objective function, the hybrid solver is not able to find the optimal solution in any run. We note that we did not adjust the Lagrange multipliers and kept the ones chosen by default by the hybrid solver. We do not rule out that customised settings for the parameters could be beneficial to the quality of the solution.
@@ 3.4 Integer/Binary Quadratic Programming (BQP)
Binary quadratic problems are NP-hard [61] and thus challenging to solve classically, even for a relatively small number of variables and constraints. As the structure of BQP resembles the Ising Hamiltonian, Equation 5, these problems are especially suited for quantum annealers [62]. By multiplying with and summing over an additional set of binary variables $x_{j}$ in Equation 9, we acquire a BQP of the form
@@
To benchmark the different solvers for this BQP, we increase the complexity by incrementally increasing the value of the constant $C$. Thereby, the complexity of the task is increased as the feasible solution space increases. This is preferred over a further increase of $N$ to avoid an unnecessarily large memory usage as the number of terms in the objective function increases with $N^{2}$.For the comparisons in this section, we choose a sufficiently large problem with $N=500$, although different sizes were tested, showing the same tendencies. The results are shown in Figure 4.
@@
Considering both solution and run time, we observe a clear computational advantage of the hybrid solver to CPLEX. We cut CPLEX off after 1000s of runtime, resulting in sub-optimal solutions when the solution space increases. Gurobi performed even worse than CPLEX for this problem and was excluded from the figure for visibility reasons. While IPOPT shows comparable solutions and run time to D-Wave, the solutions are still not optimal. D-Wave finds consistently optimal solutions in less computational time, which again results in no error bars. To more clearly visualise this, we show the relative difference in solution quality between D-Wave’s hybrid solver and the objective function value of both IPOPT and CPLEX in Figure 5.
@@
We see from Figure 5 that CPLEX performs well in the lower complexity region and that D-Wave does not find exclusively better solutions than CPLEX here; however, due to the capped computational time, D-Wave outperforms consistently for higher values of $C$. While IPOPT’s relative difference becomes increasingly small, CPLEX, with a capped solve-time of 1000s, seems to settle at a deviation of $\sim 25\%$ from the optimal solution. This shows that D-Wave’s hybrid solvers are a suitable alternative for BQPs, even showing a computational advantage for this specific case. However, in the field of optimisation, there is a limited amount of problems that correspond to these pure BQPs.

### v4

@@ 1 Introduction
In gate-based QC, the classical, binary bits are replaced by controllable qubits capable of being in non-binary states. Adiabatic quantum annealers, however, function on a different principle. They prepare an ensemble of qubits in the ground state of some initial Hamiltonian, before allowing the system to evolve abiding by the adiabatic theorem [22] to some final state encoding the desired solution. This principle will be further explained in Section˜2.1.
@@
This research aims to address this gap by utilising the largest publicly available adiabatic quantum annealers, last updated in August 2023, to solve various optimisation problems, including integer/binary linear programming (BLP) with linear and quadratic constraints, integer/binary quadratic programming (BQP), and mixed-integer linear programming. To assess performance, we systematically compare D-Wave’s Advantage quantum annealers, as described in [40], with the state-of-the-art classical solvers CPLEX [41], Gurobi [42], and IPOPT [43]. Recognising the limitations of fully QC algorithms, we employ D-Wave’s hybrid solvers [44]. This approach seamlessly integrates the strengths of both classical and quantum computation, as explained in more detail in Section˜2.5. This study serves as a valuable resource for researchers across various fields dealing with optimisation problems, enabling them to assess the suitability of current D-Wave quantum annealers for their respective applications.
The subsequent sections of the paper are organised as follows: Section˜2 provides a brief introduction to the physical and technical background of quantum annealing, encompassing the adiabatic theorem and D-Wave’s architecture in their treatment of constraints and the hybrid computing approach employed. Section˜3 presents various case studies for different classes of optimisation problems. Then, Section˜4 showcases the application of D-Wave’s hybrid constrained quadratic model (CQM) solver [45] on an energy system problem formulated as Mixed-Integer Linear Programming (MILP).
@@ 2.1 Adiabatic theorem
To perform calculations using a quantum annealer, the physical collection of qubits, hereafter referred to as the system, is prepared in an easy-to-solve initial Hamiltonian $H_{I}$. After the system has been prepared in $H_{I}$, external magnetic fields are applied to change the initial $H_{I}$ into a complex Hamiltonian $H_{F}$. Suppose this process follows the adiabatic theorem described above. In that case, the system will remain in its ground state for the duration of the anneal, and the final state will be the solution of the complex Hamiltonian. As $\sigma_{x},\sigma_{z}]\neq 0$, $H_{I}$ and $H_{F}$ do not share a common set of eigenstates. As a result, the system must undergo quantum transitions to evolve from one eigenstate to another. If the two Hamiltonians did commute, the eigenstates would remain unchanged throughout the evolution, preventing any quantum dynamics from occurring. The procedure can slowly anneal the system from the quantum Hamiltonian (Equation˜5) to the solution of the classical Ising spin system shown in Equation˜2. After the annealing process, the collapsed spin states will present a low-energy solution. The annealing quantum Ising Hamiltonian is as follows:
@@
The annealing process begins at $t=0$ and ends at time $t=t_{f}$, where $A(0)\gg B(0)$ and $A(t_{f})\ll B(t_{f})$. At time $t=t_{f}$, the qubits have been de-phased to a classical system and the Pauli matrix can be replaced by the classical $s_{i}=\pm 1$, resulting in Equation˜2.
@@ 2.2 D-Wave architecture
The most powerful QPU available is the Pegasus topology [50]. In the Pegasus topology, each qubit is coupled to 15 other qubits, which is a great improvement from the 6 couplings of the previous generation. It consists of a square lattice of unit cells which themselves contain 24 qubits. The Advantage QPU, used for the computations performed in this work, is a 16x16 lattice of unit cells, totalling $24\times 16\times(16-1)=5760$ qubits. Due to the densely packed spatial arrangement of qubits and the physical manufacturing process, it is expected that some qubits may exhibit suboptimal performance. D-Wave estimates that approximately 5% of the qubits in the Pegasus topology may have imperfections. This dense graph structure of the Pegasus systems has allowed for a significant increase in the number of useful qubits, making the Advantage QPU the most promising system to solve the optimisation problems presented in Sections˜3.1, 3.4 and 3.3.
@@ 2.3 Constraints
where $b$ is some constant that has to be met. However, when confronted with a linear inequality constraint of the form $P({\bf{x}})=\left(\sum_{i}^{N}a_{i}x_{i}-b\right)\leq 0$, auxiliary slack variables, ${\bf{y}}=\{y_{j}:\forall j\in J\}$ and their corresponding weights $w_{j}$ are used to transform inequality constraints into equality constraints [56] so that Equation˜6 can be written as
@@
with coupling values $Q$ and summing over all constraints $k$. With this approach, constrained optimisation problems can be reformulated into a solvable QUBO. We note that other approaches than Lagrangian relaxation have been considered and are actively researched to account for the constraints [57]. Furthermore, the penalty strength $\lambda$ in Equation˜8 has to be chosen with care. Contrary to classical computing, there is a far more limited dynamic range of possible pre-factors and coupling strengths that can be applied. Too large values cause a tremendous amount of distortion of the distribution in the ground state. At the same time, too small values will not ensure the satisfaction of the constraint as breaking it could then still result in optimal, unfeasible solutions. The best penalty strength is thereby the lowest possible value that still satisfies the constraints. There is no consensus on how to consistently guarantee this behaviour, and only some guidelines can be found either empirically or algebraically [58, 33].
@@ 3 Case study
This section presents different case studies and their results. The exploration covers various classes of optimisation problems: Integer/Binary Linear Programming (BLP), while also considering the influence of additional constraints of both linear and quadratic nature in Section˜3.1, and Integer/Binary Quadratic Programming (BQP) in Section˜3.4. To assess the performance of D-Wave’s Advantage system for industrial use, we benchmark the solution quality and computational time of D-Wave’s LeapHybridCQMSolver [45] against results produced by the widely adopted classical solvers CPLEX [41], Gurobi [42], and IPOPT [43] (version 3.14.12). The classical solvers are executed on a dedicated computer cluster equipped with dual 3.5 GHz Intel Xeon Gold 5115 CPUs, each featuring 10 cores and 96 GB of RAM. Noting here that for the time comparison, we measured the total run time as outputted by the cloud service, which includes the time it takes to return the full sample-set of on average 100 sample solutions. From these solutions, we first filter out the feasible set before selecting the solution with the lowest energy value. The time required for these two steps was not included in the computational time measurement for obtaining sufficient solutions. While these steps typically take less than 1s on average, their impact may only be considerable for very short total run times.
@@ 3.1 Integer/Binary Linear Programming (BLP)
where $\mu_{i}$ are random constants between (0,1) and $x_{i}\in\{0,1\}$ are binary variables. This restricts the sum over all binary variables to equal a fixed constant $C$. We first run the solvers with 10,000 to 100,000 binary variables. A comparison of the solutions and the computational time for this range of binary variables is shown in Figure˜1(a). Here, we average over 5 runs, each providing a sampleset of approximately 100 solutions from which the best feasible one is selected. As we experienced little deviation in the results we deemed higher statistics not to be necessary. Additionally, as access to quantum computers is still very limited, large statistics are not realisable. IPOPT is designed to handle large-scale nonlinear problems [43] and is not suited for integer variables. As IPOPT approximates the binary variables with continuous ones with very small bounds around zero and one, the derived objective function value with these numbers can score a lower value than what is actually true. We excluded IPOPT’s objective function values in Figure˜1(b), however, we kept this solver throughout this work as it shows important results in Section˜3.4.
@@
Looking at Figure˜1(a), D-Wave stays relatively constant in computation time of about 6s up to 38,000 variables, from whereon it increases more drastically when the number of variables is increased further. Viewing the objective function value, we see that IPOPT behaves similarly to the other solvers for a low number of variables, but starts diverging from the rest as the number of variables increases. It can also be noted that the computational time of IPOPT increases rapidly compared to the other classical solvers. CPLEX and Gurobi start below 1s and gradually increase with an increasing number of variables. D-Wave finds the same optimal solution as Gurobi and CPLEX, resulting in three overlapping lines. The observed sawtooth pattern of the solutions is due to the randomness of the $\mu$ –values as they are newly drawn for each run with a specific seed. When comparing computational time in Figure˜1(b), the hybrid solver requires significantly more time as it approaches its scalability limit. This behaviour could be explained by the need for an extremely high amount of partitioning into subproblems as the number of variables increases [35]. Although the optimal solution has been found, D-Wave’s hybrid solvers show no computational advantage.
@@ 3.2 Increasing the complexity by adding more constraints
Next, we focus on determining whether D-Wave could continue to find the optimal solution at its scalability limit when increasing the complexity of the BLP (Equation˜9) by adding additional constraints. As the constraints are added as penalty terms, we want to see if the solution quality can be maintained even though the computational time is significantly higher. We consider some arbitrary, although feasible, constraints listed below.
@@
The objective function value and computational time averaged over 5 runs, together with the minimum and maximum deviation as vertical lines, are shown in Figure˜2.
@@ 3.3 Integer/Binary Linear Programming with quadratic constraints
In the previous section, all added constraints were linear. The examination, in this section, involves the consideration of a quadratic constraint while keeping the objective function linear. To this end, we added another factor to the constraint from Equation˜10 so that it now reads
@@
where $x_{i,j}\in\{0,1\}$, resulting in $N^{2}$ terms. At first, we viewed the solution quality and computational time for an increasing number of variables, shown in Figure˜3(a), followed by a stepwise increase in the complexity of the constraint by increasing the value of $C$ in Figure˜3(b).
Figure 3: Mean objective function value (top) and computational time (bottom) comparison between two classical solvers and D-Wave’s hybrid solver for the BLP with a single quadratic constraint. In (a) the complexity scales with the number of variables. In (b), the variables are fixed to 500, whereas the complexity scales with increasing $C$ as stated in Equation˜10. All results are averaged over 5 runs, and the vertical lines show the minimum and maximum deviation from the mean value. Gurobi’s solve time is significantly larger than all other solvers and is thereby excluded here for visibility.
@@
We experience similar behaviour when stepwise increasing the value of the constraint. D-Wave is again slightly off the optimal solution, however, it stays relatively constant in computational time and solves the problem almost consistently faster than IPOPT. Although the solution is suboptimal, it can be seen in Figure˜3(b) that also for quadratic constraints with a linear objective function, the hybrid solver is not able to find the optimal solution in any run. We note that we did not adjust the Lagrange multipliers and kept the ones chosen by default by the hybrid solver. We do not rule out that customised settings for the parameters could be beneficial to the quality of the solution.
@@ 3.4 Integer/Binary Quadratic Programming (BQP)
Binary quadratic problems are NP-hard [61] and thus challenging to solve classically, even for a relatively small number of variables and constraints. As the structure of BQP resembles the Ising Hamiltonian, Equation˜5, these problems are especially suited for quantum annealers [62]. By multiplying with and summing over an additional set of binary variables $x_{j}$ in Equation˜9, we acquire a BQP of the form
@@
To benchmark the different solvers for this BQP, we increase the complexity by incrementally increasing the value of the constant $C$. Thereby, the complexity of the task is increased as the feasible solution space increases. This is preferred over a further increase of $N$ to avoid an unnecessarily large memory usage as the number of terms in the objective function increases with $N^{2}$.For the comparisons in this section, we choose a sufficiently large problem with $N=500$, although different sizes were tested, showing the same tendencies. The results are shown in Figure˜4.
@@
Considering both solution and run time, we observe a clear computational advantage of the hybrid solver to CPLEX. We cut CPLEX off after 1000s of runtime, resulting in sub-optimal solutions when the solution space increases. Gurobi performed even worse than CPLEX for this problem and was excluded from the figure for visibility reasons. While IPOPT shows comparable solutions and run time to D-Wave, the solutions are still not optimal. D-Wave finds consistently optimal solutions in less computational time, which again results in no error bars. To more clearly visualise this, we show the relative difference in solution quality between D-Wave’s hybrid solver and the objective function value of both IPOPT and CPLEX in Figure˜5.
@@
We see from Figure˜5 that CPLEX performs well in the lower complexity region and that D-Wave does not find exclusively better solutions than CPLEX here; however, due to the capped computational time, D-Wave outperforms consistently for higher values of $C$. While IPOPT’s relative difference becomes increasingly small, CPLEX, with a capped solve-time of 1000s, seems to settle at a deviation of $\sim 25\%$ from the optimal solution. This shows that D-Wave’s hybrid solvers are a suitable alternative for BQPs, even showing a computational advantage for this specific case. However, in the field of optimisation, there is a limited amount of problems that correspond to these pure BQPs.